OpenStudy (anonymous):

Find the work done by a force of 200 pounds acting in the direction -i + 2j in moving an object 75 feet from (0,0) to (-75,0). A) 13,416.1 ft-lb B)6708.2 ft-lb C) 15,000.0 ft-lb D) 8944.9 ft-lb

OpenStudy (anonymous):

I know how to do the majority of this problem, but the thing i am not sure of is what direction \[-i+2j\] is

OpenStudy (anonymous):

\[W=||F||*||AB||*\cos \theta\]

OpenStudy (anonymous):

so... \[W = 200 * 75*\cos \theta\] I am lacking the ability/knowledge to convert \[-i+2j\] into an angle

OpenStudy (amistre64):

break it into its component parts the force triangle is similar to the vectors triangle

OpenStudy (amistre64):